OpenStudy (anonymous):

Will give a medal! Please help with this log equation? Is my answer correct?

OpenStudy (anonymous):

What is the solution of log(2x − 6) 256 = 4? is it x = 5?

OpenStudy (jdoe0001):

how did you get "5"?

OpenStudy (anonymous):

I don't remember.. ._. i was playing around with a calculator and that number came up. It's one of the options.. so.. :/

OpenStudy (jdoe0001):

hmmm have you covered logarithms yet?

OpenStudy (anonymous):

The thing is that my math course is absolute crap.. so in every lesson they give you 2-3 examples and then expect you to know everything just from those examples. :/

OpenStudy (anonymous):

I've gotten the gist of it, but it's more complex problems like this that I struggle with. Could you help me?

OpenStudy (jdoe0001):

well.... lemme start off by using the exponential notation then \(\bf \large log_{\color{red}{ a}}{\color{blue}{ b}}=y\implies {\color{red}{ a}}^y={\color{blue}{ b}} \\ \quad \\ \quad \\ log_{{\color{red}{ 2x - 6}}}({\color{blue}{ 256}}) = 4\implies ({\color{red}{ 2x - 6}})^4={\color{blue}{ 256}}\) so... now... can you find a number that raised by 4 gives 256?

OpenStudy (anonymous):

4, right?

OpenStudy (jdoe0001):

yeap thus \(\bf log_{{\color{red}{ 2x - 6}}}({\color{blue}{ 256}}) = 4\implies ({\color{red}{ 2x - 6}})^4={\color{blue}{ 256}}\implies (2x-6)^4=4^4\) and you can pretty much get it from there I think

OpenStudy (jdoe0001):

2x-6=4 and then just get "x"

OpenStudy (anonymous):

thank you! I have one more that makes no sense. I've tried plugging it into my graphing calculator and into different online calculators and none of them are coming up right. Can I post it here? It's really simple, I promise.

OpenStudy (jdoe0001):

k

OpenStudy (anonymous):

Here. :/

OpenStudy (anonymous):

I tried graphing all of them all of them were too far to the right. Did I miss something? @jdoe0001

OpenStudy (anonymous):

Nevermind, got it. lol.

OpenStudy (jdoe0001):

the function is clearly an exponential one so "3" something so